OpenStudy (anonymous):

find the Vertical Asymptote; r(x) = 1/(a-(x-b)^2) Assuming a and b are >0 i got 2b and (a-b^2) as answers but they are wrong. Any help? This is for my Calculus class, and we are working on Family of Functions, if that helps.

OpenStudy (solomonzelman):

At which values of x is the denominator equal to zero?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a-(x-b)^2=0 }\) \(\large\color{black}{ \displaystyle a=(x-b)^2 }\) \(\large\color{black}{ \displaystyle \pm\sqrt{a}=x-b }\) \(\large\color{black}{ \displaystyle b\pm\sqrt{a}=x }\)

OpenStudy (anonymous):

thank you! i was expanding the equation

OpenStudy (solomonzelman):

I didn't think that finding asymptotes is calculus, but in any case I hope that helped :)